Product Z is sold to wholesalers for $5.43 and sold to end consumers for $9.99. The size of the market is $34,000,000 annually (based on retail sales); product Z’s share (in dollars) of this market is 28%. The fixed costs involved in manufacturing and managing Product Z are $1,200,000 and the variable costs involved in manufacturing and managing Product Z are $0.91 per unit. The advertising budget for Product Z is $1,000,000. Salespeople are paid entirely by an 8.5% commission based on the manufacturer’s selling price.

Strategy 2: Lowering the price

7. Calculate the total sales volume (in units) needed to maintain the current profit level if the manufacturer lowers its price by 25%.

8. Calculate the market share needed to maintain the current profit level if the manufacturer lowers its price by 25%.

Strategy 3: Increasing the sales commission

9. Calculate the total sales volume (in units) needed to maintain the current profit level if the manufacturer increases the sales force’s commission to 15%.

10. Calculate the market share needed to maintain the current profit level if the manufacturer increases the sales force’s commission to 15%
Answer #1
Per unit $
Selling price, retail = 9.99
Variable cost per unit:
manufacturing 0.91
selling commission 0.85 (9.99*8.5%)
Total Variable cost per unit 1.76
Contribution per unit 8.23
Fixed cost:
manufacturing 1200000
Advertising 1000000
Total Fixed cost 2200000
Current Profit:
Units sold 952953
Sales (28%) 9520000
Less:Variable costs: 1676387
Contribution 7843613
Less:Fixed costs 2200000
Current Profit: 5643613
7)
If selling price reduced by 25% to = 9.99*0.75 = 7.49
Total sales volume (in units):
Revised Contribution per unit (A)= 7.49 - 1.76 = 5.73
Contribution required (B) = $7843613
Units sold = B / A = 7843613/5.73 = 1368868 units
Sales value = 1368868 * 7.49 = $10252821
8) New market share = 10252821/34000000 = 30.16%
9) If commission raised to 15%.
New Variable cost = 0.91 + (9.99*15%) = $2.41 per unit
Total sales volume (in units):
Revised Contribution per unit (A)= 9.99 - 2.41 = 7.58
Contribution required (B) = $7843613
Units sold = B / A = 7843613/7.58 = 1034777 units
Sales value = 1034777 * 9.99 = $10337422
10) New market share = 10337422/34000000 = 30.40%